OpenStudy (anonymous):

Find vertical asymtotes for (x^3-10x^2+16x)/(x^2-8x) I know that x cannot equal 0 or 8, where do I go from here?

OpenStudy (anonymous):

ok I gotcha

OpenStudy (anonymous):

ya still there?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

one sec while I figure it out

OpenStudy (anonymous):

is there any way you can plug that in using the equation editor for me?

OpenStudy (anonymous):

\[(x^3-10x^2+16x)/(x^2-8x)\]

OpenStudy (anonymous):

Tell me the steps on how to attain the answer, I don't need to know the answer I need to know the how rather than end, thanks!

OpenStudy (anonymous):

This problem is sort of a trick question.\[\frac{x^3-10 x^2+16 x}{x^2-8 x}=x-2 \]A plot is attached.

OpenStudy (anonymous):

could I see this done arithmetically?

OpenStudy (anonymous):

there are no vertical asymptotes

OpenStudy (anonymous):

\[\frac{x^3-10 x^2+16 x}{x^2-8 x}=\frac{((x-8) (x-2) x)}{((x-8) x)}=x-2 \]

OpenStudy (anonymous):

am I right rob?

OpenStudy (anonymous):

@llkramer Yes.